# Energy Conditions and Quantum Information

### v1

@@ 4.1 Black hole entropy and holography
The path to holography begins with the thermal nature of black holes. Hawking’s discovery Hawking:1974rv ; Hawking:1975vcx that black holes emit blackbody radiation implies that they possess a well-defined temperature. From the laws of black hole mechanics, first formulated by Bekenstein Bekenstein:1973ur and Hawking, a consistent thermodynamic description emerged:
@@
$S_{\rm BH}$ $=\frac{A}{4G\hbar}\quad\mbox{(Bekenstein-Hawking entropy)}$ (91)
where $T_{\rm H}$ is called the Hawking temperature given by the surface gravity $\kappa$, and $S_{\rm BH}$ called the Bekenstein-Hawking entropy formula, proportional to the area $A$ of a cross-section of the event horizon as seen in Sec. 2.6.
@@
Since black holes are believed to be the densest and most entropic objects in nature, they offer a natural upper bound on the information content in a given region. This inspired the conjecture that the fundamental theory underlying quantum gravity is holographic: {itembox}[l]Conjecture 4.1.1 (Holographic principle) The complete description of a gravitational system in a bulk spacetime is equivalently formulated as a lower-dimensional, non-gravitational quantum field theory living on its boundary. A concrete and precise realization of this principle was proposed by Maldacena Maldacena:1997re in 1997, now famously known as the AdS/CFT correspondence.
@@
Figure 15: [Left] The cylinder depicts a conformally compactified AdS spacetime. The cylinder’s interior corresponds to the AdS bulk $M$ and its surface represents the conformal boundary $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. The diamond-shaped subregion on $\partial M$ corresponds to $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{4},\eta_{\mu\nu})$. This subregion is enclosed by the red curves, which represent the future and past null infinity ($\mathscr{I}^{\pm}\subset\partial M$), and a single point $i^{0}\in\partial M$, which represents spatial infinity. [Right] 22-dimensional section of the conformal diagram of AdS spacetime. Two vertical double lines represent conformal infinity (i.e., the surface of the cylinder). The triangular subregion inside $M$ is covered by the horospherical chart (also known as the Poincare chart) given by (94), where $\infty>z>0$ inside $M$. Each vertical hyperboloid represents a constant-$z$ hypersurface, which is a $d$-dimensional Minkowski spacetime with the metric $(\ell/z)^{2}\eta_{\mu\nu}$. The boundary is located at $z=0$.
The AdS metric $g_{MN}$ in $(d+1)$-dimension can be expressed in the horospherical (Poincare) chart as
@@
where $\ell$ denotes the AdS curvature radius defined by $\ell^{2}=-d(d-1)/2\Lambda$ and where $z\rightarrow 0$ corresponds to infinity. Multiplying $g_{MN}$ by ${z^{2}}/\ell^{2}$ and taking the limit $z\rightarrow 0$, we obtain the $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{d},\eta_{\mu\nu})$ as part of the conformal boundary $\partial M$, which is considered as the arena for a quantum field theory (see Figure 15). As (part of) the $d$-dimensional boundary spacetime $(\partial M,g_{\mu\nu})$, one can also consider different geometries, such as $d$-dimensional de Sitter, AdS, or the static Einstein universe ${\mathbb{R}}\times S^{d-1}$ (see the metric (32) for $d=4$ case), other than the Minkowski spacetime. As for the $(d+1)$-dimensional bulk geometry $(M,g_{MN})$, one can, in principle, consider any asymptotically AdS spacetime, such as those including black holes inside but at large distances approaching the above AdS metric. In general, the AdS/CFT correspondence asserts the following: {itembox}[l]Conjecture 4.1.2 (AdS/CFT correspondence) A duality between a $(d+1)$-dimensional string theory defined in asymptotically Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) without gravity residing on its conformal boundary.
@@
where the logarithmic term appears only when $d$ is even, and $g^{(2k+1)}_{\mu\nu}(x)=0$ for any integer $k$ satisfying $0\leqslant 2k+1<d$. This is called the Fefferman-Graham (FG) expansion. For given $g^{(0)}_{\mu\nu}(x)$, all the subleading coefficients $g^{(n)}_{\mu\nu}(x)$ with $n<d$ are determined solely by the boundary metric $g^{(0)}_{\mu\nu}(x)$ and the curvature tensors with respect to $g^{(0)}_{\mu\nu}(x)$, while the coefficient $g^{(d)}_{\mu\nu}(x)$ is related to the expectation value of the boundary field stress-energy tensor $\langle T_{\mu\nu}\rangledeHaro$:2000vlm : {itembox}[l]Formula 4.1.3 (Holographic stress-energy tensor) The renormalized stress-energy tensor on the boundary quantum field is given in terms of the FG expansion coefficient as,
$\langle T_{\mu\nu}\rangle=\dfrac{d\ell^{d-1}}{16\pi G_{d+1}}g^{(d)}_{\mu\nu}(x)+{X}_{\mu\nu}\,.$ (97)
Here $X_{\mu\nu}$ represents gravitational anomaly, which vanishes when the boundary dimension $d$ is odd. This formula stems from the fact that the AdS/CFT correspondence essentially equates the partition functions—and thus their effective actions—of the bulk and boundary theories. It is generally a formidable task to derive the left-hand side by using conventional quantum field theoretic methods. In the above formula, this task is considerably simplified by the use of the AdS/CFT correspondence, which replaces it with the FG expansion of the classical metric in the AdS bulk spacetime. In Sec. 4.4 and 4.5, we will discuss how the holographic stress-tensor derived from this formula can be used to examine the ANEC for strongly coupled quantum fields. However, before that, we will discuss more on recent progress in the holographic principle and understanding the black hole information paradox in the next two subsections.
@@ 4.2 AdS/CFT correspondence and the black hole information paradox
[…]
@@ 4.3 Entanglement entropy, the Ryu-Takayanagi formula, and the Island paradigm
A pivotal insight in the effort to resolve the black hole information paradox is that entropy, particularly entanglement entropy, plays a fundamental role in the structure of quantum gravity. In semiclassical gravity, the entropy of a black hole is famously captured by the Bekenstein-Hawking formula, {itembox}[l]Formula 4.3.1 (Bekenstein-Hawking entroy)
$S_{\text{BH}}=\frac{A}{4G\hbar}\,$ (98)
where $A$ denotes the area of the event horizon as mentioned at the beginning of this section. This area-law scaling stands in sharp contrast to the volume-law behavior characteristic of entropy in local quantum field theory, hinting at a radical departure from conventional locality. It suggests that the underlying degrees of freedom in quantum gravity must be organized in a nonlocal fashion.
This conceptual leap gained concrete form in the context of the AdS/CFT correspondence, especially with the introduction of the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef . For a spatial subregion $A$ of the boundary CFT, the RT formula computes its entanglement entropy as {itembox}[l]Formula 4.3.2 (Ryu-Takayanagi formula)
$S_{A}=\min_{\gamma}\left[\frac{\mathrm{Area}(\gamma_{A})}{4G}\right]\,$, (99)
where $\gamma_{A}$ is the codimension-2 minimal surface in the bulk AdS geometry, anchored to $\partial A$ and homologous to $A$. This prescription geometrizes entanglement: it relates a quantum information-theoretic quantity to a geometric property of spacetime, suggesting that entanglement is not merely a diagnostic tool but a foundational principle underlying gravitational physics.
@@
$S_{A}=S_{\text{BH}}=\frac{A_{\text{horizon}}}{4G}.$ (100)
@@
These bring us to the notion of quantum extremal surfaces (QES), developed by Faulkner, Lewkowycz, and Maldacena Faulkner:2013ana and further formalized by Engelhardt and Wall Engelhardt:2014gca . The QES prescription replaces the classical minimal-area surface with a surface that extremizes the generalized entropy: {itembox}[l]Formula 4.3.3 (Quantum extremal surface)
$S_{\text{gen}}(\gamma)=\mbox{Ext}_{\gamma}\left[\frac{\mathrm{Area}(\gamma)}{4G}+S_{\text{bulk}}(\Sigma_{\gamma})\right]\,$, (101)
where $\Sigma_{\gamma}$ is the bulk region bounded by $\gamma$ and the boundary region $A$, and $S_{\text{bulk}}(\Sigma_{\gamma})$ is the von Neumann entropy of quantum fields in that region. The entanglement entropy of region $A$ is then computed by extremizing this generalized entropy over all candidate surfaces $\gamma$ anchored to $\partial A$. Note that the idea of including quantum corrections from matter fields in this way is conceptually rooted in Bekenstein’s notion of generalized entropy Bekenstein:1974ax . There, Bekenstein proposed that the total entropy should consist of both the area term and the matter entropy outside the horizon, anticipating the structure of $S_{\text{gen}}$ long before the advent of the AdS/CFT correspondence, even before the discovery of Hawking radiation.
@@
A major conceptual advance came with the development of the island formula, especially in the context of evaporating black holes coupled to an external non-gravitating bath. This framework was pioneered by Almheiri, Engelhardt, Marolf, and Maxfield Almheiri:2019hni and Penington Penington:2019npb , and further developed by Almheiri:2019psf . In such a setup, the entanglement entropy $S_{R}$ of the Hawking radiation $R$ is given by {itembox}[l]Formula 4.3.4 (Island formula)
$S_{R}=\mbox{Ext}_{I}\left[\frac{\mathrm{Area}(\partial I)}{4G}+S_{\text{bulk}}(R\cup I)\right]\,.$ (102)
Here $I$ denotes a candidate island region inside the black hole, and $\partial I$ is its boundary (see Figure 18). At early times, the minimal configuration has no island, reproducing Hawking’s prediction of monotonically increasing radiation entropy. However, after the Page time, a new configuration appears where a nontrivial island emerges near the black hole horizon that becomes dominant, yielding a smaller generalized entropy. As a result of this island, the entropy of the radiation begins to decrease after the Page time, successfully reproducing the Page curve.
@@
Moreover, this flow of insights is not unidirectional. While boundary entanglement patterns determine the bulk geometry, the consistency and causal structure of the bulk spacetime also impose powerful constraints on the boundary theory. A notable example is the derivation of boundary energy conditions, such as the averaged null energy condition (ANEC), from bulk causality and modular flow. In the next section, we explore how this bidirectional logic further deepens our understanding of quantum gravity.
@@ 4.4 Causal constraints and the ANEC in holography
where $k^{\mu}$ is a future-directed null vector and the expectation value is given by the holographic Formula 4.1.1 deHaro:2000vlm . The key insight in the Kelly-Wall argument is that any violation of the ANEC would imply a breakdown of boundary causality via its holographic encoding in bulk geometry. More precisely, if bulk signals could propagate in a manner that violates the causal structure of the boundary theory, then the duality itself would become inconsistent. To formalize this, they consider null deformations of boundary subregions and analyze the corresponding changes in the bulk extremal surfaces computing their entanglement entropy. A violation of the ANEC would induce deformations that allow extremal surfaces to access regions outside the causal domain of the boundary, thereby violating entanglement wedge nesting and the monotonicity of relative entropy.
This perspective is closely related to a foundational result by Gao and Wald Gao:2000ga , who showed that in asymptotically AdS spacetimes satisfying the NEC, bulk causal propagation respects boundary causal structure. More precisely, consider an asymptotically AdS spacetime $(M,g_{MN})$ and its conformal boundary $\partial M$. Suppose a pair of boundary two points $p,q\in\partial M$ are connected by a null geodesic $\gamma$ which is lying entirely in $\partial M$ and is achronal with respect to the boundary metric $g^{(0)}_{\mu\nu}$. If there is a timelike curve through the bulk $M$ which also connects the boundary two points $p$ and $q$, then the entire spacetime $M\cup\partial M$ is said to admit a bulk-shortcut. {itembox}[l]Theorem 4.4.1 (Gao-Wald theorem) Consider an asymptotically AdS spacetime $M$ which has no causal pathology, such as a naked singularity or causality violation, in the bulk. Then, under the (A)NEC and the null generic condition in the bulk, there exists no bulk-shortcut and the fastest possible causal curve connecting two boundary points $p,q$ is an achronal null geodesic $\gamma$ lying entirely on the boundary $\partial M$.

### v2

@@ 4.1 Black hole entropy and holography
The path to holography begins with the thermal nature of black holes. Hawking’s discovery Hawking:1974rv ; Hawking:1975vcx that black holes emit blackbody radiation implies that they possess a well-defined temperature. Building on Bekenstein’s insight that horizon area encodes entropy Bekenstein:1973ur , the four laws of black hole mechanics were formulated by Bardeen, Carter, and Hawking Bardeen:1973gs , yielding a consistent thermodynamic description:
@@
which is called the Hawking temperature given by the surface gravity $\kappa$, and
  Formula 4.1.1 (Bekenstein-Hawking entropy)   $S_{\text{BH}}=\frac{k_{\rm B}A}{4G\hbar}\,.$ (93)     This is called the Bekenstein-Hawking entropy formula, proportional to the area $A$ of a cross-section of the event horizon as seen in Sec. 2.6.
@@
Since black holes are believed to be the densest and most entropic objects in nature, they offer a natural upper bound on the information content in a given region. This inspired the conjecture that the fundamental theory underlying quantum gravity is holographic:
  Conjecture 4.1.2 (Holographic principle)   The complete description of a gravitational system in a bulk spacetime is equivalently formulated as a lower-dimensional, non-gravitational quantum field theory living on its boundary.     A concrete and precise realization of this principle was proposed by Maldacena Maldacena:1997re in 1997, now famously known as the AdS/CFT correspondence.
@@
Figure 15: [Left] The cylinder depicts a conformally compactified AdS spacetime. The cylinder’s interior corresponds to the AdS bulk $M$ and its surface represents the conformal boundary $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. The diamond-shaped subregion on $\partial M$ corresponds to $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{d},\eta_{\mu\nu})$. This subregion is enclosed by the red curves, which represent the future and past null infinity ($\mathscr{I}^{\pm}\subset\partial M$), and a single point $i^{0}\in\partial M$, which represents spatial infinity. [Right] 22-dimensional section of the conformal diagram of AdS spacetime. Two vertical double lines represent conformal infinity (i.e., the surface of the cylinder). The triangular subregion inside $M$ is covered by the horospherical chart (also known as the Poincaré chart) given by (96), where $\infty>z>0$ inside $M$. Each vertical hyperboloid represents a constant-$z$ hypersurface, which is a $d$-dimensional Minkowski spacetime with the metric $(\ell/z)^{2}\eta_{\mu\nu}$. The boundary is located at $z=0$.
The AdS metric $g_{MN}$ in $(d+1)$-dimension can be expressed in the horospherical (Poincaré) chart as
@@
where $\ell$ denotes the AdS curvature radius defined by $\ell^{2}=-d(d-1)/2\Lambda$ and where $z\rightarrow 0$ corresponds to infinity. Multiplying $g_{MN}$ by ${z^{2}}/\ell^{2}$ and taking the limit $z\rightarrow 0$, we obtain the $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{d},\eta_{\mu\nu})$ as part of the conformal boundary $\partial M$, which is considered as the arena for a quantum field theory (see Figure 15). As (part of) the $d$-dimensional boundary spacetime $(\partial M,g_{\mu\nu})$, one can also consider different geometries, such as $d$-dimensional de Sitter, AdS, or the static Einstein universe ${\mathbb{R}}\times S^{d-1}$ (see the metric (32) for $d=4$ case), other than the Minkowski spacetime.
As for the $(d+1)$-dimensional bulk geometry $(M,g_{MN})$ in the AdS/CFT correspondence, one can, in principle, consider any asymptotically AdS spacetime, such as those including black holes inside but at large distances approaching the above AdS metric. In general, the AdS/CFT correspondence asserts the following:
  Conjecture 4.1.3 (AdS/CFT correspondence)   A duality between a $(d+1)$-dimensional string theory defined in asymptotically Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) without gravity residing on its conformal boundary.
@@
where the logarithmic term appears only when $d$ is even, and $g^{(2k+1)}_{\mu\nu}(x)=0$ for any integer $k$ satisfying $0\leqslant 2k+1<d$. This is called the Fefferman-Graham (FG) expansion. For given $g^{(0)}_{\mu\nu}(x)$, all the subleading coefficients $g^{(n)}_{\mu\nu}(x)$ with $n<d$ are determined solely by the boundary metric $g^{(0)}_{\mu\nu}(x)$ and the curvature tensors with respect to $g^{(0)}_{\mu\nu}(x)$, while the coefficient $g^{(d)}_{\mu\nu}(x)$ is related to the expectation value of the boundary field stress-energy tensor $\langle T_{\mu\nu}\rangledeHaro$:2000vlm :
  Formula 4.1.4 (Holographic stress-energy tensor)   The renormalized stress-energy tensor on the boundary quantum field is given in terms of the FG expansion coefficient as, $\langle T_{\mu\nu}\rangle=\dfrac{d\ell^{d-1}}{16\pi G_{d+1}}g^{(d)}_{\mu\nu}(x)+{X}_{\mu\nu}\,.$ (99)     Here $X_{\mu\nu}$ represents gravitational anomaly, which vanishes when the boundary dimension $d$ is odd. This formula stems from the fact that the AdS/CFT correspondence essentially equates the partition functions—and thus their effective actions—of the bulk and boundary theories. It is generally a formidable task to derive the left-hand side by using conventional quantum field theoretic methods. In the above formula, this task is considerably simplified by the use of the AdS/CFT correspondence, which replaces it with the FG expansion of the classical metric in the AdS bulk spacetime. In Sec. 4.4 and 4.5, we will discuss how the holographic stress-tensor derived from this formula can be used to examine the ANEC for strongly coupled quantum fields. However, before that, we will discuss more on recent progress in the holographic principle and understanding the black hole information paradox in the next two subsections.
@@ 4.2 AdS/CFT correspondence and the black hole information paradox
Figure 16: An entangled pair of particles A and B in a spherically symmetric spacetime of black hole formation. The thin dotted hyperboloids depict curves of constant area radius $r$ and $\mathscr{H}$$^{+}$ represents the event horizon. Particle A falls into the black hole and moves toward smaller $r$, whereas B moves toward larger $r$. As a result, the distance between the two particles increases with time. Particle B eventually escapes to $\mathscr{I}^{+}$ as Hawking radiation. This process generates entanglement between the black hole interior and the outgoing Hawking radiation.
@@ 4.3 Entanglement entropy, the Ryu-Takayanagi formula, and the Island paradigm
A pivotal insight in the effort to resolve the black hole information paradox is that entropy, particularly entanglement entropy, plays a fundamental role in the structure of quantum gravity. In semiclassical gravity, the entropy of a black hole is famously captured by the Bekenstein-Hawking formula (93). This area-law scaling stands in sharp contrast to the volume-law behavior characteristic of entropy in local quantum field theory, hinting at a radical departure from conventional locality. It suggests that the underlying degrees of freedom in quantum gravity must be organized in a nonlocal fashion.
This conceptual leap gained concrete form in the context of the AdS/CFT correspondence, especially with the introduction of the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef . For a spatial subregion $A$ of the boundary CFT, the RT formula computes its entanglement entropy as [^16] In the Ryu–Takayanagi formula, the entanglement entropy has a UV divergence of the form $S_{A}\propto\mathrm{Area}(\gamma_{A})/\epsilon^{2}$. Here $\epsilon$ is a short-distance cutoff in the boundary CFT: in AdS Poincaré coordinates, instead of extending the minimal surface all the way to the boundary at $z=0$, one terminates it at $z=\epsilon$, which regulates the divergence.
  Formula 4.3.1 (Ryu-Takayanagi formula)   $S_{A}=\min_{\gamma_{A}}\left[\dfrac{k_{\rm B}\mathrm{Area}(\gamma_{A})}{4G\hbar}\right]\,.$ (100)     Here $\gamma_{A}$ is the codimension-2 minimal surface in the bulk AdS geometry, anchored to $\partial A$ and homologous to $A$. This prescription geometrizes entanglement: it relates a quantum information-theoretic quantity to a geometric property of spacetime, suggesting that entanglement is not merely a diagnostic tool but a foundational principle underlying gravitational physics.
@@
$S_{A}=S_{\text{BH}}=\dfrac{k_{\rm B}A_{\text{horizon}}}{4G\hbar}\,.$ (101)
@@
These bring us to the notion of quantum extremal surfaces (QES), developed by Faulkner, Lewkowycz, and Maldacena Faulkner:2013ana and further formalized by Engelhardt and Wall Engelhardt:2014gca . The QES prescription replaces the classical minimal-area surface with a surface that extremizes the generalized entropy:
$S_{\mathrm{gen}}(\gamma)=\dfrac{k_{\rm B}\mathrm{Area}(\gamma)}{4G\hbar}+S_{\mathrm{bulk}}(\Sigma_{\gamma})\,$, (102)
where $\Sigma_{\gamma}$ is the bulk region on an achronal slice bounded by $\gamma$ and the boundary region $A$, and $S_{\text{bulk}}(\Sigma_{\gamma})$ is the von Neumann entropy of quantum fields in that region. The quantum extremal surface $\gamma_{A}$ is then defined as follows:
  Formula 4.3.2 (Quantum extremal surface)   $S(A)=\mbox{min-ext}_{\gamma}S_{\mathrm{gen}}(\gamma).$ (103)     Here $\mbox{min-ext}_{\gamma}$ represents the following process: first search for extremal surface(s) $\gamma_{A}$ (homologous to $A$) that satisfy the extremality $\delta_{\gamma_{A}}S_{\mathrm{gen}}=0$, and then choose the one which attains the minimum among all such extremal surfaces if there are many such $\gamma_{A}$. The surface obtained by this min-ext prescription (among QES homologous to $A$) provides the correct entanglement entropy $S(A)$ Engelhardt:2014gca . Note that the inclusion of matter entanglement in this way is conceptually rooted in Bekenstein’s generalized entropy Bekenstein:1974ax . There, Bekenstein proposed that the total entropy should consist of both the area term and the matter entropy outside the horizon, anticipating the structure of $S_{\text{gen}}$ long before the advent of the AdS/CFT correspondence, even before the discovery of Hawking radiation.
@@
A major conceptual advance came with the development of the island formula, especially in the context of evaporating black holes coupled to an external non-gravitating bath Almheiri:2019hni . In practice, this means working in a regime where gravity is dynamical only in the black hole region, while the bath lives on a fixed background. This framework was pioneered by Almheiri, Engelhardt, Marolf, and Maxfield Almheiri:2019hni and Penington Penington:2019npb , and further developed by Almheiri:2019psf . In such a setup, the entanglement entropy $S_{R}$ of the Hawking radiation $R$ is given by
  Formula 4.3.3 (Island formula)   $S_{R}=\mbox{min-ext}_{I}\left[\frac{k_{\rm B}\mathrm{Area}(\partial I)}{4G}+S_{\text{bulk}}(R\cup I)\right]\,.$ (104)     Here $I$ denotes a candidate island region inside the black hole, and $\partial I$ is its boundary (see Figure 18). At early times, the minimal configuration has no island, reproducing Hawking’s prediction of monotonically increasing radiation entropy. However, after the Page time, a new configuration appears where a nontrivial island emerges near the black hole horizon that becomes dominant, yielding a smaller generalized entropy. As a result of this island, the entropy of the radiation begins to decrease after the Page time, successfully reproducing the Page curve [^17] In this argument, we assume the existence of a nice foliation of regular (partial) Cauchy surfaces which cover the entire process, ranging from black hole formation to the final stage (an immediate vicinity) of black hole evaporation. For asymptotically flat spacetimes, each of these time slices can be anchored at some boundary surface in the asymptotic region, sometime referred to as cutoff surface or anchor cruve Penington:2019npb . The calculation of QES and the determination of the Page time can be performed based on such time slices. .
@@
Moreover, this flow of insights is not unidirectional. While boundary entanglement patterns determine the bulk geometry, the consistency and causal structure of the bulk spacetime also impose powerful constraints on the boundary theory. A notable example is the derivation of boundary energy conditions, such as the averaged null energy condition (ANEC), from bulk causality. In the next section, we explore how this bidirectional logic further deepens our understanding of quantum gravity.
@@ 4.4 Causal constraints and the ANEC in holography
where $k^{\mu}$ is a future-directed null vector and the expectation value is given by the holographic Formula 4.1.4 deHaro:2000vlm . The key insight in the Kelly-Wall argument is that any violation of the ANEC would imply a breakdown of boundary causality via its holographic encoding in bulk geometry. More precisely, if bulk signals could propagate in a manner that violates the causal structure of the boundary theory, then the duality itself would become inconsistent. To formalize this, they consider null deformations of boundary subregions and analyze the corresponding changes in the bulk extremal surfaces computing their entanglement entropy. A violation of the ANEC would induce deformations that allow extremal surfaces to access regions outside the causal domain of the boundary, thereby violating entanglement wedge nesting and the monotonicity of relative entropy.
This perspective is closely related to a foundational result by Gao and Wald Gao:2000ga , who showed that in asymptotically AdS spacetimes satisfying the NEC, bulk causal propagation respects boundary causal structure. More precisely, consider an asymptotically AdS spacetime $(M,g_{MN})$ and its conformal boundary $\partial M$. Suppose a pair of boundary points $p,q\in\partial M$ are connected by a null geodesic $\gamma$ which is lying entirely in $\partial M$ and is achronal with respect to the boundary metric $g^{(0)}_{\mu\nu}$. If there is a timelike curve through the bulk $M$ which also connects the boundary two points $p$ and $q$, then the entire spacetime $M\cup\partial M$ is said to admit a bulk-shortcut.
  Theorem 4.4.1 (Gao-Wald theorem)   Consider an asymptotically AdS spacetime $M$ which has no causal pathology, such as a naked singularity or causality violation, in the bulk. Then, under the (A)NEC and the null generic condition in the bulk, there exists no bulk-shortcut and the fastest possible causal curve connecting two boundary points $p,q$ is an achronal null geodesic $\gamma$ lying entirely on the boundary $\partial M$.